In a microwave oven, water molecules are forced to
Answer #1

multiplied with two the fomula is E.p(1-cos(theta)) cos(180) = -1

so

the resultant becomes 2.E.p

From the formula given,

E = p.E

=6.2* 10-30 x 2000

=1.24 x 10-26 J

This is for one OH bond so resultant is 2* 1.24 * 10^-26 = 2.5 * 10^-26J